Is [ (1,1), (2,1), (2,2), (1,2)] a function?
Is f(x)= 3x-3 a function?
Evaluate f(a+1) for f defined in f(x) = 3x-3.
What is the domain and range for f in f(x) = 3x-3.

Answer by robertb(5830)

{(1,1), (2,1), (2,2), (1,2)} is not a function, because f(1) = 1 and 2, f(2) = 1 and 2.
f(x) is a function, because it is linear.
f(a+1) = 3(a+1) - 3 = 3a + 3 - 3 = 3a.
Domain and range of f(x) are both the set of all real numbers.